OpenStudy (anonymous):

Classify the triangle based on side lengths 25, 15, 9.

Mehek (mehek14):

15^2 + 9^2 = 25^2 225 + 81 = 625 ^ not equal so it is not a right angle triangle

Mehek (mehek14):

are there any answer choices?

OpenStudy (anonymous):

@Mehek14 right acute obtuse no triangle can be formed with given side lengths

Mehek (mehek14):

ok so it not right

OpenStudy (anonymous):

obtuse or no triangle..

Mehek (mehek14):

ok lets try by drawing it

Mehek (mehek14):

it is an scalene triangle because it has three different sides and pythagorean theorem is not proved in it so it is not a right angled triangle so it is obtuse